Hi. I am having trouble with this problem:
Expressed in a + bi form, 5/ 3 + i is equuivalent to...
A) 15 - 5i, B) (3/2) - (1/2i), C) (5/3) - 5i, D) (15/8) - (5/8i)

Answer by nerdybill(7384)

Answer is B
.
Multiply by (3-i)/(3-i):
5/(3+i) * (3-i)/(3-i)
= 5(3-i)/(3+i)(3-i)
= 5(3-i)/(9-3i+3i+1)
= 5(3-i)/(9+1)
= 5(3-i)/10
= (15-5i)/10
= 15/10 - 5i/10
= (3/2) - (1/2)i